it says find the common solution(s).
y=3x^2
y=3x+6


4y + 3x=12
2x+3y=1

Answer by stanbon(75887)

find the common solution(s).
y=3x^2
y=3x+6
-----
3x^2 = 3x + 6
x^2 = x + 2
x^2-x-2 = 0
Factor:
(x-2)(x+1) = 0
x = 2 or x = -1
---
If x = 2, y = 12
If x = -1, y = 3
========================

4y + 3x=12
2x+3y=1
-----
Rearrange:
3x + 4y = 12
2x + 3y = 1
-----
Multiply top equation by 2
Multiply bottom equation by 3
----
6x + 8y = 24
6x + 9y = 3
------
Subtract to solve for "y":
y = -21
---
Solve for "x" using 2x + 3y = 1
---
2x +3*-21 = 1
2x -63 = 1
2x = 64
x = 32
=============
Cheers,
Stan H.
===============